A grocer mixes cashews and almonds. A mixture of 4 pounds of cashews and 6 pounds of almonds cost $62. A mixture of 2 pounds of cashews and 8 pounds of almonds cost $54. What is the cost per pound of cashew and almonds?
Graphically?

Answer by richwmiller(17219)

4c+6a=62
2c+8a=54
multiply by 2
4c+16a=108
4c+6a=62
subtract
10a= 46
a=$4.60 per lb
2c+8*4.6=54
2c=17.2
c=$8.60 per lb